find the area of a regular octagon with a perimeter of 96 centimeters
Answer by anantha(86)

sol:
octagon having 8 sides
let side be x
perimeter=x+x+x+x+x+x+x+x=8x
according to the problem perimeter=96 centimeters
8x=96
x=96/8=12 centimeters
side=12 cm
area of the octagon=8*area of equiliteral triangle
area of an equiliteral triangle=root3/4 *(side^2)
=root3/4*(12^2)
root3/4*(12*12)
=root3*3*12=36*root3
now area of octagon=8*36*root3=288 *root3 cm^2